# Stability of discrete-symmetry flocks: sandwich state, traveling domains and motility-induced pinning

### v1

@@ Appendix B Sandwich formation in the active Ising model (AIM) at reduced transverse diffusion
Figure 13: (color online) Formation of the sandwich state in the AIM for small transverse diffusion. (a) $D_{\perp}=0.1$ and (b) $D_{\perp}=0.3$. Full reversal of the initial liquid phase is observed for larger values of $D_{\perp}$, (c) $D_{\perp}=0.5$, and (d) $D_{\perp}=1$. Colorbar represents per-site magnetization $m_{i}=n_{i}^{+}-n_{i}^{-}$, where $n_{i}^{+}$ and $n_{i}^{-}$ are the number of particles with $s=+1$ and $s=-1$ at site $i$, respectively. Parameters: $D_{\parallel}=1$, $\beta=2$, $v=1$, $\rho_{0}=10$, $r_{d}=10$, $\rho_{0}^{d}=5\rho_{0}$, $L_{x}=1000$, and $L_{y}=100$.
@@
We simulate the AIM (as defined in Ref. Benvegnen et al. (2023)) on a two-dimensional lattice where each particle carries an Ising spin variable, $s=\pm 1$, indicating its self-propulsion direction. We decompose the diffusion constant $D$ into two components, $D_{\perp}$ (along ±$\pm y$ direction) and $D_{\parallel}$ (along ±$\pm x$ direction). Particles hop to one of their four neighbors at diffusion rate $2\left(D_{\perp}+D_{\parallel}\right)$, self-propel to a neighboring site on the right $(s=+1)$ or left $(s=-1)$ at rate $v$, and can flip their spin state $(s\to-s)$ at rate $\gamma e^{-\beta sm_{i}/\rho_{i}}$, where $m_{i}=n_{i}^{+}-n_{i}^{-}$ and $\rho_{i}=n_{i}^{+}+n_{i}^{-}$ are the local magnetization and density at site $i$, respectively. We set $\gamma=1$ and the Monte Carlo time unit of our simulation is taken as $\Delta t=\left(2D_{\perp}+2D_{\parallel}+v+e^{\beta}\right)^{-1}$. The artificial droplet excitation follows the protocol described in Sec. III.2.
@@
The initial system is prepared by inserting a droplet of $s=+1$ particles in the high-density liquid phase of $s=-1$ particles. We vary $D_{\perp}$, keeping $D_{\parallel}=v=1$. Fig. 13 shows that a sandwich state can also emerge in the AIM (previously unreported) when $D_{\perp}$ is low. For $D_{\perp}=0.1$ and $D_{\perp}=0.3$ [Fig. 13(a–b)], the droplet cannot expand along the transverse (±$(\pm y)$ direction due to limited diffusion and does not form a comet-like structure, as shown in Ref. Benvegnen et al. (2023), instead, a sandwich state of $s=+1$ (red) and $s=-1$ (blue) particles appears. As $D_{\perp}$ increases [Fig. 13(c–d)], the comet-like structure of the droplet reappears due to enhanced transverse diffusion, resulting in a complete reversal of the initial liquid phase, as reported in Ref. Benvegnen et al. (2023).
@@
As demonstrated in Fig. 13, the outcome of an artificial droplet excitation, whether it leads to a sandwich state or a full reversal of the initial liquid phase, depends on the transverse diffusion $D_{\perp}$, where the biased movement of the droplet is confined in $1d$ (±$\pm x$ direction). The transverse length of the system $L_{y}$ also plays a crucial role, as a complete phase reversal requires the growing droplet front to connect the boundaries along the transverse direction ($y=1$ and $y=L_{y}$), which is only possible if sufficient transverse diffusion is present. For $D_{\perp}=1$ (maximum value), the droplet front will inevitably connect the transverse boundaries and trigger a full reversal, regardless of $L_{y}$. However, when $D_{\perp}$ is small, a full reversal can only occur if $L_{y}$ is also small. Therefore, the critical $D_{\perp}$, above which a full reversal takes place, increases with $L_{y}$. For instance, in Fig. 13(b), if $L_{y}=40$ is considered instead of $L_{y}=100$, a full reversal is observed rather than the sandwich state. However, if $L_{y}=40$ is maintained and $D_{\perp}$ is reduced, the sandwich state reappears.
@@ Appendix C Impact of spatial anisotropy on the emergence of the APM sandwich state
Appendix D Counter-propagating droplet dynamics for varying $\beta$
@@
In Sec. III.2, we examined the effect of artificial droplet excitation deep within the ordered phase of the APM by considering large $\beta=1$ and density $\rho_{0}=10$. We found that, under these conditions, counter-propagating droplet excitation leads to a steady-state characterized by a partially reversed sandwich configuration, in the absence of any lattice anisotropy. In Fig. 15(a), we explore how this behavior changes when $\beta$ is reduced, thereby enhancing thermal fluctuations in a square geometry ($L_{x}=L_{y}=200$). For $\beta>0.8$, we find that the system consistently exhibits the sandwich state across all values of $\epsilon$, with the droplet-induced lane becoming narrower as $\epsilon$ increases. For $\beta=0.75$ and $\epsilon=1.5$ [Fig. 15(a)], the counter-propagating droplet is observed to completely replace the initial ordered phase. This full reversal occurs within a central range of $\epsilon$ values $(1.2\leq\epsilon\leq 1.8)$, near the reorientation transition Chatterjee et al. (2020); Mangeat et al. (2020), which itself depends on $\beta$. Outside this range, the system predominantly exhibits a sandwich state. Fig. 15(b) shows the time-evolution of the system with transverse droplet excitation at $\beta=1$ and $\epsilon=1.5$ where the steady-state is characterized by mutually orthogonal motion of two high-density clusters of the droplet (green) and the initial liquid state (blue). This mutual orthogonality in motion, reinforced by strong directional bias and alignment interactions, allows the two high-density clusters to stably coexist without coalescing, even over long times.
Appendix E Metastability in the APM with a different hopping mechanism
@@
Figure 16: (color online) Spontaneous nucleation in the APM with a different hopping mechanism. (a) Time evolution snapshots of the modified APM (see Appendix E) with initially ordered configuration, $\sigma=4$ (black domain), exhibit spontaneous nucleation of droplets which indicate fragility of the LRO phase similar to Fig. 8. (b–d) Steady-state snapshots with increasing $v_{0}$ show reduced spontaneous nucleation of droplets and formation of longitudinally moving bands (signifying liquid-gas coexistence region), which characterizes high-velocity particle regimes. Parameters: $D_{0}=0.05$, $\beta=1$, $\rho_{0}=5$, and $L_{x}=L_{y}=1024$. The four different colors in the color bar represent the different spin states.
@@
Inspired by the AIM model proposed in Ref. Woo and Noh (2024), here we propose a moderately different version of the APM and show that, similar to the AIM, this model also exhibits spontaneous nucleation of droplets and metastability of the LRO phase. In this model, we change the hopping mechanism by keeping the flipping rate unchanged [Eq. (3)]. The particles now hop to one of their four neighbors at a rate of $4D_{0}$ and self-propel to a neighboring site according to spin state $\sigma$ at a rate $v_{0}$. Therefore, the total hopping rate becomes $4D_{0}+v_{0}$ ($4D$ in the previous model). This hopping is equivalent to Eq. (4) with $D_{0}=D(1-\epsilon/3)$ and $v_{0}=4D\epsilon/3$. Since $\epsilon$ is constrained within the range $0\leqslant\epsilon\leqslant 3$, the velocity $v_{0}=4D_{0}\epsilon/(3-\epsilon)$ in this model is theoretically unbounded (at $\epsilon=3$, $v_{0}\to\infty$). The Monte Carlo time unit is also adjusted and is now given by $\Delta t=\left[4D_{0}+v_{0}+\exp(4\beta)\right]^{-1}$ with $4D_{0}+v_{0}=4D$.
Fig. 16 shows the time-evolution [Fig. 16(a)] and steady-state snapshots [Fig. 16(b–d)] for various values of self-propulsion velocity $v_{0}$. Fig. 16(a) shows that the high-density liquid phase is also not stable under the new hopping mechanism with diffusion $D_{0}=0.05$ and $v_{0}=1$ $(\epsilon=2.5)$ akin to the AIM Woo and Noh (2024). The liquid band, composed of particles in the $\sigma=4$ state (black domain), initially moves downward. Subsequently, spontaneous nucleation of clusters predominantly composed of $\sigma=1$ (red) and $\sigma=3$ (blue) state particles, moving right and left, respectively, emerges within high-density regions ($t=131$). This nucleation disrupts the long-range order (LRO), leading to a new steady state characterized by short-range ordered (SRO) clusters of particles from all states ($t=311$ and $t=3571$). Due to the high velocity, the clusters move longitudinally without significant lateral diffusion. This behavior persists at $v_{0}=2$ $(\epsilon\simeq 2.73)$ [Fig. 16(b)], but for $v_{0}=3$ $(\epsilon\simeq 2.81)$ [Fig. 16(c)] and $v_{0}=4$ $(\epsilon\simeq 2.86)$[Fig. 16(d)], steady longitudinal bands form. This is because increasing $v_{0}$ increases the total hopping rate, which in the original APM depended only on $D$ and was independent of $\epsilon$ [see Eq. (4)]. A liquid phase normally arises at large $\beta$, because the state flipping rate is higher, and for spontaneous nucleation to occur, the total hopping rate must be small relative to $W_{\rm flip}$, which depends solely on $\beta$ and local density. As $v_{0}$ increases, both the biased and total hopping rates increase, stabilizing the band motion characteristic of the coexistence region. Therefore, it can be concluded that as the total hopping rate becomes comparable to the flipping rate, spontaneous nucleation ceases to exist. This explains why spontaneous nucleation was not observed for $D=1$ in the original APM Chatterjee et al. (2020); Mangeat et al. (2020), since varying $\epsilon$ does not affect the total hopping rate, which remains fixed at $4D$. Spontaneous nucleation in the APM with the original hopping rule can only be observed by reducing the diffusion constant $D$, as shown in Fig. 8.
Appendix F Droplet pinning at low temperature
@@
Figure 17: (color online) Droplet jamming at small temperatures. (a) Snapshots at different times, $t$, showing droplet pinning in the 4-state APM. A droplet of $\sigma=1$ particles, initially placed in an ordered region of $\sigma=3$ particles, becomes pinned over time, highlighting the pinning effect under low diffusion and high $\beta$. A movie (movie11) of the same can be found at Ref. Chatterjee et al. (2025). (b) Magnetization profile at the interface between the droplet state (red solid line) and the ordered phase (blue solid line) at $t=160$, showing a pronounced peak that indicates a jammed, high-magnetization region at the droplet front. (c) Position of the droplet interface front ($x_{f}$) relative to the center, showing saturation over time, indicating the formation of a jammed, pinned interface. Parameters: $D=0.3$, $\beta=2$, $\epsilon=2.5$, $\rho_{0}=10$, $r_{d}=6$, $\rho_{0}^{d}=5\rho_{0}$, $L_{x}=200$, $L_{y}=50$.

### v2

@@ Appendix A Temporal evolution of a counter-propagating droplet
Figure 13: (color online) Formation of the sandwich state in the AIM for small transverse diffusion. (a) $D_{\perp}=0.1$ and (b) $D_{\perp}=0.3$. Full reversal of the initial liquid phase is observed for larger values of $D_{\perp}$, (c) $D_{\perp}=0.5$, and (d) $D_{\perp}=1$. Colorbar represents per-site magnetization $m_{i}=n_{i}^{+}-n_{i}^{-}$, where $n_{i}^{+}$ and $n_{i}^{-}$ are the number of particles with $s=+1$ and s=1s=-1 at site $i$, respectively. Parameters: $D_{\parallel}=1$, $\beta=2$, $v=1$, $\rho_{0}=10$, $r_{d}=10$, $\rho_{0}^{d}=5\rho_{0}$, $L_{x}=1000$, and $L_{y}=100$.
@@ Appendix B Sandwich formation in the active Ising model (AIM) at reduced transverse diffusion
We simulate the AIM (as defined in Ref. benvegnen2023meta) on a two-dimensional lattice where each particle carries an Ising spin variable, $s=\pm 1$, indicating its self-propulsion direction. We decompose the diffusion constant $D$ into two components, $D_{\perp}$ (along $\pm y$ direction) and $D_{\parallel}$ (along $\pm x$ direction). Particles hop to one of their four neighbors at diffusion rate $2\left(D_{\perp}+D_{\parallel}\right)$, self-propel to a neighboring site on the right $(s=+1)$ or left (s=1)(s=-1) at rate $v$, and can flip their spin state $(s\to-s)$ at rate $\gamma e^{-\beta sm_{i}/\rho_{i}}$, where $m_{i}=n_{i}^{+}-n_{i}^{-}$ and $\rho_{i}=n_{i}^{+}+n_{i}^{-}$ are the local magnetization and density at site $i$, respectively. We set $\gamma=1$ and the Monte Carlo time unit of our simulation is taken as $\Delta t=\left(2D_{\perp}+2D_{\parallel}+v+e^{\beta}\right)^{-1}$. The artificial droplet excitation follows the protocol described in Sec. III.2.
@@
The initial system is prepared by inserting a droplet of $s=+1$ particles in the high-density liquid phase of s=1s=-1 particles. We vary $D_{\perp}$, keeping $D_{\parallel}=v=1$. Fig. 13 shows that a sandwich state can also emerge in the AIM (previously unreported) when $D_{\perp}$ is low. For $D_{\perp}=0.1$ and $D_{\perp}=0.3$ [Fig. 13(a–b)], the droplet cannot expand along the transverse $(\pm y)$ direction due to limited diffusion and does not form a comet-like structure, as shown in Ref. benvegnen2023meta, instead, a sandwich state of $s=+1$ (red) and s=1s=-1 (blue) particles appears. As $D_{\perp}$ increases [Fig. 13(c–d)], the comet-like structure of the droplet reappears due to enhanced transverse diffusion, resulting in a complete reversal of the initial liquid phase, as reported in Ref. benvegnen2023meta.
@@
As demonstrated in Fig. 13, the outcome of an artificial droplet excitation, whether it leads to a sandwich state or a full reversal of the initial liquid phase, depends on the transverse diffusion $D_{\perp}$, where the biased movement of the droplet is confined in $1d$ ($\pm x$ direction). The transverse length of the system $L_{y}$ also plays a crucial role, as a complete phase reversal requires the growing droplet front to connect the boundaries along the transverse direction ($y=1$ and $y=L_{y}$), which is only possible if sufficient transverse diffusion is present. For $D_{\perp}=1$ (maximum value), the droplet front will inevitably connect the transverse boundaries and trigger a full reversal, regardless of $L_{y}$. However, when $D_{\perp}$ is small, a full reversal can only occur if $L_{y}$ is also small. Therefore, the critical $D_{\perp}$, above which a full reversal takes place, increases with $L_{y}$. For instance, in Fig. 13(b), if $L_{y}=40$ is considered instead of $L_{y}=100$, a full reversal is observed rather than the sandwich state. However, if $L_{y}=40$ is maintained and $D_{\perp}$ is reduced, the sandwich state reappears.
@@ Appendix C Impact of spatial anisotropy on the emergence of the APM sandwich state
Appendix D Profiles and stability of the sandwich state
Figure 15: (color online) Density and magnetization profiles of the sandwich state. Time-averaged (a) density and (b) magnetization profiles of the sandwich state for increasing droplet radius $r_{d}$. Parameters: $D=1$, $\beta=1$, $\epsilon=2.7$, $\rho_{0}=10$, and $\rho_{0}^{d}=5\rho_{0}$. $L_{x}=200$, $L_{y}=50$.
Fig. 15 shows the time-averaged density and magnetization profiles, integrated over the x-coordinate, of the sandwich state formed by droplets of $\sigma=1$ embedded in a $\sigma=3$ liquid, for various droplet radii $r_{d}$. For all $r_{d}$, the density $\overline{\rho}(y)$ is almost uniform (except near the interfaces) around the average density after introducing the droplet: $\rho_{\rm tot}=(N+\Delta N)/L_{x}L_{y}$, whereas the magnetization $\overline{m_{\sigma=1}}(y)$ peaks at the middle, where the droplet have been introduced, and becomes negative outside where the magnetization $\overline{m_{\sigma=3}}(y)$ dominates. Therefore, the sandwich state is not density-segregated, and only the internal polarization varies along the transverse direction.
Figure 16: (color online) Stability of the sandwich state. (a) Hydrodynamic theory for $D=1$, $\beta=0.9$, $\rho_{0}=3$, $L_{x}=200$, and $L_{y}=50$. (b) Numerical simulations for $D=1$, $\beta=1$, $\rho_{0}=10$, $L_{x}=200$, and $L_{y}=50$. At fixed $\epsilon$, the sandwich state is stable if the width $w_{\rm sw}$ is between $w_{\min}$ and $w_{\max}$. For a prepared sandwich, widths below $w_{\min}$ collapse to a uniform single-polarity liquid matching the outer domains, while widths above $w_{\max}$ expand and convert the entire system to a liquid phase of inner polarity.
Fig. 16 summarizes the stability of the sandwich state by solving the hydrodynamic equations, Eq. (5) [Fig. 16(a)]; and from numerical simulations [Fig. 16(b)]. For fixed $\epsilon$, we analyze the stability a sandwich state with initial width $w_{\rm sw}$, where the rectangular domain is separated into two vertically-organized lanes of state $\sigma=1$ and $\sigma=3$ of size $w_{\rm sw}$ and $L_{y}-w_{\rm sw}$, respectively, with equal density. The sandwich state remains stable only if the initial width is between two thresholds: a minimal width $w_{\min}(\epsilon)$ and a maximal width $w_{\max}(\epsilon)$. If $w_{\rm sw}<w_{\min}$, the system goes to a liquid of state $\sigma=3$, and if $w_{\rm sw}>w_{\max}$, the system goes to a liquid of state $\sigma=1$. With symmetry arguments, we can identify the relation $w_{\max}=L_{y}-w_{\min}$. The hydrodynamic and numerical results agree qualitatively on this stability analysis.
Appendix E Counter-propagating droplet dynamics for varying $\beta$
@@
In Sec. III.2, we examined the effect of artificial droplet excitation deep within the ordered phase of the APM by considering large $\beta=1$ and density $\rho_{0}=10$. We found that, under these conditions, counter-propagating droplet excitation leads to a steady-state characterized by a partially reversed sandwich configuration, in the absence of any lattice anisotropy. In Fig. 17(a), we explore how this behavior changes when $\beta$ is reduced, thereby enhancing thermal fluctuations in a square geometry ($L_{x}=L_{y}=200$). For $\beta>0.8$, we find that the system consistently exhibits the sandwich state across all values of $\epsilon$, with the droplet-induced lane becoming narrower as $\epsilon$ increases. For $\beta=0.75$ and $\epsilon=1.5$ [Fig. 17(a)], the counter-propagating droplet is observed to completely replace the initial ordered phase. This full reversal occurs within a central range of $\epsilon$ values $(1.2\leq\epsilon\leq 1.8)$, near the reorientation transition chatterjee2020flocking; mangeat2020flocking, which itself depends on $\beta$. Outside this range, the system predominantly exhibits a sandwich state. Fig. 17(b) shows the time-evolution of the system with transverse droplet excitation at $\beta=1$ and $\epsilon=1.5$ where the steady-state is characterized by mutually orthogonal motion of two high-density clusters of the droplet (green) and the initial liquid state (blue). This mutual orthogonality in motion, reinforced by strong directional bias and alignment interactions, allows the two high-density clusters to stably coexist without coalescing, even over long times.
Appendix F Metastability in the APM with a different hopping mechanism
@@
Figure 18: (color online) Spontaneous nucleation in the APM with a different hopping mechanism. (a) Time evolution snapshots of the modified APM (see Appendix F) with initially ordered configuration, $\sigma=4$ (black domain), exhibit spontaneous nucleation of droplets which indicate fragility of the LRO phase similar to Fig. 8. (b–d) Steady-state snapshots with increasing $v_{0}$ show reduced spontaneous nucleation of droplets and formation of longitudinally moving bands (signifying liquid-gas coexistence region), which characterizes high-velocity particle regimes. Parameters: $D_{0}=0.05$, $\beta=1$, $\rho_{0}=5$, and $L_{x}=L_{y}=1024$. The four different colors in the color bar represent the different spin states.
@@
Inspired by the AIM model proposed in Ref. jd2024MIP, here we propose a moderately different version of the APM and show that, similar to the AIM, this model also exhibits spontaneous nucleation of droplets and metastability of the LRO phase. In this model, we change the hopping mechanism by keeping the flipping rate unchanged [Eq. (3)]. The particles now hop to one of their four neighbors at a rate of $4D_{0}$ and self-propel to a neighboring site according to spin state $\sigma$ at a rate $v_{0}$. Therefore, the total hopping rate becomes $4D_{0}+v_{0}$ ($4D$ in the previous model). This hopping is equivalent to Eq. (4) with $D_{0}=D(1-\epsilon/3)$ and $v_{0}=4D\epsilon/3$. Since $\epsilon$ is constrained within the range $0\leqslant\epsilon\leqslant 3$, the velocity $v_{0}=4D_{0}\epsilon/(3-\epsilon)$ in this model is theoretically unbounded (at $\epsilon=3$, $v_{0}\to\infty$). The Monte Carlo time unit is also adjusted and is now given by $\Delta t=\left[4D_{0}+v_{0}+\exp(4\beta)\right]^{-1}$ with $4D_{0}+v_{0}=4D$.
@@
Figure 19: (color online) Droplet jamming at small temperatures. (a) Snapshots at different times, $t$, showing droplet pinning in the 4-state APM. A droplet of $\sigma=1$ particles, initially placed in an ordered region of $\sigma=3$ particles, becomes pinned over time, highlighting the pinning effect under low diffusion and high $\beta$. A movie (movie11) of the same can be found at Ref. zenodo. (b) Magnetization profile at the interface between the droplet state (red solid line) and the ordered phase (blue solid line) at $t=160$, showing a pronounced peak that indicates a jammed, high-magnetization region at the droplet front. (c) Position of the droplet interface front ($x_{f}$) relative to the center, showing saturation over time, indicating the formation of a jammed, pinned interface. Parameters: $D=0.3$, $\beta=2$, $\epsilon=2.5$, $\rho_{0}=10$, $r_{d}=6$, $\rho_{0}^{d}=5\rho_{0}$, $L_{x}=200$, $L_{y}=50$.
Fig. 18 shows the time-evolution [Fig. 18(a)] and steady-state snapshots [Fig. 18(b–d)] for various values of self-propulsion velocity $v_{0}$. Fig. 18(a) shows that the high-density liquid phase is also not stable under the new hopping mechanism with diffusion $D_{0}=0.05$ and $v_{0}=1$ $(\epsilon=2.5)$ akin to the AIM jd2024MIP. The liquid band, composed of particles in the $\sigma=4$ state (black domain), initially moves downward. Subsequently, spontaneous nucleation of clusters predominantly composed of $\sigma=1$ (red) and $\sigma=3$ (blue) state particles, moving right and left, respectively, emerges within high-density regions ($t=131$). This nucleation disrupts the long-range order (LRO), leading to a new steady state characterized by short-range ordered (SRO) clusters of particles from all states ($t=311$ and $t=3571$). Due to the high velocity, the clusters move longitudinally without significant lateral diffusion. This behavior persists at $v_{0}=2$ $(\epsilon\simeq 2.73)$ [Fig. 18(b)], but for $v_{0}=3$ $(\epsilon\simeq 2.81)$ [Fig. 18(c)] and $v_{0}=4$ $(\epsilon\simeq 2.86)$[Fig. 18(d)], steady longitudinal bands form. This is because increasing $v_{0}$ increases the total hopping rate, which in the original APM depended only on $D$ and was independent of $\epsilon$ [see Eq. (4)]. A liquid phase normally arises at large $\beta$, because the state flipping rate is higher, and for spontaneous nucleation to occur, the total hopping rate must be small relative to $W_{\rm flip}$, which depends solely on $\beta$ and local density. As $v_{0}$ increases, both the biased and total hopping rates increase, stabilizing the band motion characteristic of the coexistence region. Therefore, it can be concluded that as the total hopping rate becomes comparable to the flipping rate, spontaneous nucleation ceases to exist. This explains why spontaneous nucleation was not observed for $D=1$ in the original APM chatterjee2020flocking; mangeat2020flocking, since varying $\epsilon$ does not affect the total hopping rate, which remains fixed at $4D$. Spontaneous nucleation in the APM with the original hopping rule can only be observed by reducing the diffusion constant $D$, as shown in Fig. 8.
Appendix G Droplet pinning at low temperature